OpenStudy (anonymous):

in 2006, there were 12,440 starbucks stores worldwide. this is approximately 1126% more stores than 10 years earlier. how many starbucks stores were there in 1996? the answer is 1015 but idk how they got that answer

OpenStudy (anonymous):

See, the equation goes this way: no. of stores in 2006 = no. of stores in 1996 + (1126/100)*no. of stores in 1996 let the no. of stores in 1996 = x So, 12440 = x+(1126/100)*x (For our convenience, let's multiply with 100 on both sides) 12440*100=100*x+(1126/100)*x*100 1244000=100x+1126x 1244000=1226x x=1244000/1226 \[x \approx 1015\]